# Adversarially-Robust Gossip Algorithms for Approximate Quantile and Mean Computations

### v2

@@ Abstract
Gossip algorithms distribute information in a scalable and efficient way by having random pairs of nodes exchange small messages. Value aggregation problems are of particular interest in this setting as they occur frequently in practice and many elegant algorithms have been proposed for computing aggregates and statistics such as averages and quantiles. An important and well-studied advantage of gossip algorithms is their robustness to message delays, network churn, and unreliable message transmissions. These crucial robustness guarantees however only hold if all nodes follow the protocol and no messages are corrupted.
In this paper, we remedy this by providing a framework to model both adversarial participants and message corruptions in gossip-style communications by allowing an adversary to control a small fraction of the nodes or corrupt messages arbitrarily. Despite this very powerful and general corruption model, we show that one can design robust gossip algorithms for many important aggregation problems. Our algorithms guarantee that almost all nodes converge to an approximately correct answer with optimal efficiency and essentially as fast as without corruptions.
@@ 1 Introduction
Distributed computing provides a resource-efficient framework for scalable algorithms that can handle the vast amounts of data produced by the increasingly interconnected and digitalized world. Designing fast and reliable algorithms to compute aggregates, such as averages, quantiles, maxima or minima, is one active area of research in this field. These algorithms are especially vital in the context of sensor networks and P2P networks [GK04, MFHH03, PS04, SBAS04].
@@
A promising class of distributed algorithms are gossip algorithms, which are typically simple, fast, scalable and robust to errors. Owing to this, many gossip algorithms have been developed for aggregation of information in the real world. Their communication protocol is strikingly simple. In these algorithms, communication between nodes occurs in (synchronous) rounds - in each round, every node selects a communication partner uniformly at random from the set of all nodes, and chooses to either push or pull a message from its partner. These messages are usually small – $O(\log n)$ bits, where $n$ is the total number of nodes.
@@
Push- and pull-based protocols each have their own unique advantages and shortcomings. In [KDG03], the authors gave a simple, fast gossip algorithm to compute the mean, and built on this to compute other aggregates, including arbitrary quantiles. The main algorithm to compute the mean is a simple yet elegant push-based protocol where each node pushes half of its current value to a random node in every round. Since such a push-protocol preserves the total sum of the values, it is not hard to see why this algorithm computes the mean, and it can be shown that it converges in $O(\log n)$ rounds with high probability. Improved algorithms that reduce the total number of messages sent were given in [CP12] and [KDN+06]. In [HMS18], the authors gave a faster, elegant and in fact optimal algorithm to compute exact quantiles in $O(\log n)$ rounds with high probability, even in the presence of random faults. Their main algorithm is a pull-based tournament-style algorithm where each node repeatedly updates its value to the median of the values of three randomly chosen nodes. This computes an approximate median extremely quickly - in $O(\log\log n+\log\frac{1}{\varepsilon})$ rounds for an $\varepsilon$-approximation - and can be bootstrapped to compute exact quantiles in $O(\log n)$ rounds.
@@
One obstacle to the real-world deployment of these algorithms is the presence of message corruptions - be they malicious or benign. In the case of some algorithms, such as the median finding algorithm in [HMS18], they are shown to be robust to random failures, i.e. where nodes fail to perform their (pull- or push-) operation. The much more challenging problem however lies in the design of algorithms that are also robust to - potentially adversarial - message corruption. This arguably also greatly increases their real-world applicability. Indeed, some algorithms break completely in the presence of even a few adversarial faults in each round. This is the case for the push-protocol in [KDG03], where if even a single node reports a value that is abnormally large, all nodes would converge to an average that is much larger than the true average. Moreover, proof techniques used there such as the mass-preservation completely break down in the presence of adversarial corruption. However, a pull-based approach like the one used in [HMS18] naturally limits the number of corrupted messages, making it a powerful paradigm for addressing adversarial message faults. Our algorithms additionally ensure that every node sends the same value in a given round, regardless of how many nodes pull from it in that round. This approach simplifies dealing with adversarial corruption and keeps the algorithm itself simpler.
@@
To ensure our algorithms are suitable for realistic settings, we strive to model faults in the most general manner possible. Some fundamental questions that arise include the amount of (computational) power that a potential adversary may have have or how many messages may be affected by corruptions. We need to strike a balance and allow enough power to retain a meaningful adversary, but not so much that it stifles all reasonable attempts at node communication and eliminates nontrivial results. In the following, we briefly sketch the setting, concentrating for convenience on the example of an adversary, but the model readily covers non-malicious message corruptions.
@@
We propose an adversary that can influence a $\beta$ fraction of the nodes in each round, which we call a $\beta$-strong adversary, see section 1.2 for a more formal definition. The adversary can corrupt any message pulled from the nodes under its influence in this round, and can arbitrarily change their contents. We do not make any assumptions on the computational power of the adversary. This model covers the special case where the same $\beta n$ nodes are controlled by the adversary across all rounds (in other words, the case where there are $\beta n$ adversarial nodes).
@@
With the introduction of such adversarial faults, it is not reasonable to ask that the nodes output exact answers. Thus, the best that we can hope for is an approximate solution, which is exactly the type of problems we investigate. In the $\varepsilon$-approximate quantile problem, the goal is to compute a value whose rank is in the interval $[(\phi-\varepsilon)n,(\phi+\varepsilon)n]$. We will start by tackling the special case $\phi=1/2$ of computing an $\varepsilon$-approximate median, and show afterwards how the problem of computing an approximation of arbitrary quantiles can be reduced to that case. We then turn to the problem of finding the (approximate) mean, which is not as robust to outliers as the median. Thus, computing the mean - even approximately - in the presence of adversarial faults is difficult, as they can sabotage any algorithm by injecting extreme outliers into the computation. The best one can hope for this $\varepsilon$-approximate mean problem is therefore to find a value that is at most an additive $\varepsilon M$ away from the true mean, where the values are assumed to be restricted to the interval $[0,M]$.
@@
The scenario with $\beta n$ adversarial nodes tells us that we cannot hope for an approximation factor $\varepsilon$ smaller than $\beta$. Indeed, one could simply have all the adversarial nodes reliably report an extremely large or an extremely small value, which would shift the quantiles of the “good nodes” by $\beta$, and the same scenario can be applied to the mean approximation problem as well. Hence, all our results have $\varepsilon\geq\beta$ as an (implicit) assumption. In some applications – and given that a $\beta$ fraction of nodes might never output a correct answer anyway – it might be acceptable that some other nodes also hold an incorrect answer, provided this number is small enough, especially if this accelerates the algorithm in return. To investigate the possibility of trading off speed with correctness, we analyze our algorithms from the perspective of parametrized correctness, which is new in this setting. Our gossip algorithms are said to have correctness level $\gamma$ if at most $\gamma n$ nodes are storing an incorrect value (e.g. a value with rank outside $[1/2-\varepsilon,1/2+\varepsilon]$ for the $\varepsilon$-approximate median problem) upon termination. Note that we can always assume that that $\gamma\geq\frac{1}{2n}$, since $\gamma<\frac{1}{n}$ already implies that there must be no incorrect node.
Paper Organization
@@
The remainder of the paper is organized as follows. We give an overview of our results in section 1.1, followed by a description of our model and an intuitive description of our three algorithms - complemented by their pseudocode - in section 1.2. Section 1.3 contains a brief discussion of related work. Then, we proceed to prove our main theorems in order: section 2 is dedicated to our median algorithm and the proof of Theorem 1; section 3 concerns itself with our quantile shifting algorithm and the proof of Theorem 2. In section 4, we prove Theorem 3, which contains the theoretical guarantees for our mean algorithm. Section 5 contains the proof of our lower bounds, including Theorem 4. Finally, in the appendix, we collect standard concentration results and prove some technical inequalities which are used throughout the paper.
1.1 Results
@@
In the rest of the paper, we say that an event $\mathcal{E}$ occurs with high probability (w.h.p.) if $\mathbb{P}\left[\mathcal{E}\right]=1-n^{-\Omega(1)}$. Our first main result gives a fast, robust algorithm for approximate median computation. The pseudocode can be found in Algorithm 1 and is complemented by the runtime and correctness guarantees provided in our first theorem.
@@
We remark that the approximation here is weaker as it depends on the promised bound on $M$. The reason for this is simple - the adversary can always claim to have a value that is very large, and this would inevitably skew the computed average to a very large value. Thus, the best one can hope for is to give an error bound as a function of a bound on the possible communicated values.
Note that in the special case where all the nodes hold values in $\{0,1\}$, the above yields an approximate counting algorithm with an additive error of at most $\varepsilon n$.
@@
We also extend the $\Omega\big{(}\log\log n+\log\frac{1}{\varepsilon}\big{)}$ lower bound given in [HMS18] to an $\Omega\big{(}\log\log\frac{1}{\gamma}+\log\frac{1}{\varepsilon}\big{)}$ lower bound in our setting, see Proposition 31 in section 5. These results prove that our median and quantile algorithms have optimal round complexity and the mean algorithm has an almost optimal round complexity. We remark that our theorems require varying upper bounds on $\beta$ in terms of $\varepsilon$. As mentioned in the introduction, it is not possible to hope for $\varepsilon$-approximation with $\beta>\varepsilon$, as in particular all adversarial nodes might always output values that are completely out of the range of the correct values. Our analysis here is likely not tight and we did not optimize these bounds since our algorithms already cover a substantial range of values. The same holds for the upper bound on $\varepsilon$ in Theorem 2. For lower bounds on $\varepsilon$, note that any $\varepsilon<\frac{1}{2n}$ would yield an exact algorithm, which is out of reach due to the presence of the adversary. Hence a lower bound which is inversely polynomial in $n$ follows naturally, where again we did not optimize for the precise exponent, and simply stated what was required for our analysis in order to prove concentration of the expected behavior of the algorithms. The key reason why we do not view this restriction as major however is a different one: For many applications, approximation factors that are a small constant fraction are completely sufficient. Moreover, this is also the parameter range where our algorithms are extremely fast. More concretely, when both $\varepsilon$ and $\gamma$ are constant, our algorithms are guaranteed to terminate in constant time, no matter how large $n$ is.
1.2 Model Definition and Algorithm Overview
@@
We consider a network of $n$ nodes that are all connected to each other (one can think of it as a complete graph on $n$ vertices). For all the problems that we study, each node is initially assigned a value that is polynomially bounded in $n$. Communication occurs in synchronous rounds, during each of which every node contacts some node chosen uniformly at random (including itself), and then either pushes or pulls a message - of size $O(\log n)$ bits - from the chosen node. As mentioned before, all our algorithms use the pull-operation only, and the message sent out by a pulled node is the same, no matter which (and how many) nodes are pulling from it.
@@
In every round, the messages sent out by $\beta n$ nodes can become corrupted. More precisely, we assume that these corruptions can be adversarial in the following sense: at the beginning of each round, the adversary can chose at most $\beta n$ nodes so that any message pulled from or pushed to these nodes can be entirely and arbitrarily modified by the adversary, i.e., the message can be changed to whatever the adversary wants. Crucially, the adversary does not know beforehand the random choice of partners of the nodes in this round. We do not assume any additional restrictions on the adversary. For example, the adversary may have unbounded computational power and know the entire history of the algorithm at any given moment of time.
@@
Before describing our new adversarially-robust algorithms, we briefly describe the state-of-the-art in the non-adversarial setting. The approximate median finding algorithm given by [HMS18] uses a pull-based tournament-style protocol. Here, every node repeatedly updates its value to the median of three randomly chosen node values. The effect is that after $O\big{(}\log\log n+\log\frac{1}{\varepsilon}\big{)}$ rounds, every node has communicated directly or indirectly with $\log n/\varepsilon^{2}$ other nodes. Notice that the problem becomes straightforward with large message sizes, as every node can collect all the information about every node it has directly or indirectly heard from, and $\log n/\varepsilon^{2}$ many uniformly randomly chosen node values are sufficient to prove concentration of the median value for this node. As the authors show, we do not need to store all such information and we can make do with just maintaining the median. Their algorithm then concludes with a majority vote over a constant number of nodes to allow for a union bound.
@@
Our algorithms retain the simplicity of these previous algorithms with minor changes, while being robust to adversarial corruptions. Some of the crucial changes include giving a more nuanced termination condition which takes into account $\beta$ and $\gamma$, and deriving tailored bounds on binomial terms to bound the effect of the adversaries on each round. We also introduce an extended second phase – now collecting values from $O\big{(}\frac{\log(1/\gamma)}{\log(1/\beta)}\big{)}$ many nodes – to ensure that at most a $\gamma$ fraction of the nodes are storing an incorrect value upon termination, as desired.
@@
More specifically, for the approximate median computation, as outlined above, we first let nodes conduct a 3-tournament, that is, in the first phase, each node pulls three node values uniformly at random and updates its own value to the median of the three. In the second, additional, phase, the algorithm then uniformly samples a number of nodes and outputs their median. The pseudocode for our approximate median algorithm is given in Algorithm 1.

### v3

@@ Abstract
Gossip algorithms distribute information in a scalable and efficient way by having random pairs of nodes exchange small messages. Value aggregation problems are of particular interest in this setting, as they occur frequently in practice, and many elegant algorithms have been proposed for computing aggregates and statistics such as averages and quantiles. An important and well-studied advantage of gossip algorithms is their robustness to message delays, network churn, and unreliable message transmissions. However, these crucial robustness guarantees only hold if all nodes follow the protocol and no messages are corrupted.
In this paper, we remedy this by providing a framework to model both adversarial participants and message corruptions in gossip-style communications by allowing an adversary to control a small fraction of the nodes or corrupt messages arbitrarily. Despite this very powerful and general corruption model, we show that robust gossip algorithms can be designed for many important aggregation problems. Our algorithms guarantee that almost all nodes converge to an approximately correct answer with optimal efficiency and essentially as fast as without corruptions.
@@ 1 Introduction
Distributed computing provides a resource-efficient framework for scalable algorithms that can handle the vast amounts of data produced by the increasingly interconnected and digitalized world. Designing fast and reliable algorithms to compute aggregates, such as averages, quantiles, maxima, or minima, is one active area of research in this field. These algorithms are especially vital in the context of sensor networks and P2P networks [GK04, MFHH03, PS04, SBAS04].
@@
A promising class of distributed algorithms are gossip algorithms, which are typically simple, fast, scalable, and robust to errors. Due to this, many gossip algorithms have been developed for the aggregation of information in the real world. Their communication protocol is strikingly simple. In these algorithms, communication between nodes occurs in (synchronous) rounds - in each round, every node selects a communication partner uniformly at random from the set of all nodes and chooses to either push or pull a message from its partner. These messages are usually small – $O(\log n)$ bits, where $n$ is the total number of nodes.
@@
Push-based and pull-based protocols each have their own unique advantages and shortcomings. In [KDG03], the authors gave a simple, fast gossip algorithm to compute the mean and built on this to compute other aggregates, including arbitrary quantiles. The main algorithm to compute the mean is a simple yet elegant push-based protocol where each node pushes half of its current value to a random node in every round. Since such a push-protocol preserves the total sum of the values, it is not hard to see why this algorithm computes the mean, and it can be shown that it converges in $O(\log n)$ rounds with high probability. Improved algorithms that reduce the total number of messages sent were given in [CP12] and [KDN+06]. In [HMS18], the authors gave a faster, elegant, and in fact optimal algorithm to compute exact quantiles in $O(\log n)$ rounds with high probability, even in the presence of random faults. Their main algorithm is a pull-based tournament-style algorithm, where each node repeatedly updates its value to the median of the values of three randomly chosen nodes. This computes an approximate median extremely quickly - in $O(\log\log n+\log\frac{1}{\varepsilon})$ rounds for an $\varepsilon$-approximation - and can be bootstrapped to compute exact quantiles in $O(\log n)$ rounds.
@@
One obstacle to the real-world deployment of these algorithms is the presence of message corruptions, be they malicious or benign. In the case of some algorithms, such as the median finding algorithm in [HMS18], they are shown to be robust to random failures, that is, where nodes fail to perform their (pull- or push-) operation. The much more challenging problem, however, lies in the design of algorithms that are also robust to - potentially adversarial - message corruption. This arguably also greatly increases their applicability in the real world. Indeed, some algorithms break completely in the presence of even a few adversarial faults in each round. This is the case for the push-protocol in [KDG03], where if even a single node reports a value that is abnormally large, all nodes would converge to an average that is much larger than the true average. Moreover, the proof techniques used there, such as the mass preservation property, completely break down in the presence of adversarial corruption. However, a pull-based approach like the one used in [HMS18] naturally limits the number of corrupted messages, making it a powerful paradigm to address adversarial message faults. Our algorithms additionally ensure that every node sends the same value in a given round, regardless of how many nodes pull from it in that round. This approach simplifies the problem of adversarial corruption and makes the algorithm itself simpler.
@@
To ensure that our algorithms are suitable for realistic settings, we strive to model faults in the most general manner possible. Some fundamental questions that arise include the amount of (computational) power that a potential adversary may have or how many messages may be affected by corruptions. We need to strike a balance and allow enough power to retain a meaningful adversary, but not so much that it stifles all reasonable attempts at node communication and eliminates nontrivial results. In the following, we briefly sketch the setting, concentrating for convenience on the example of an adversary, but the model readily covers nonmalicious message corruptions.
@@
We propose an adversary that can influence a $\beta$ fraction of the nodes in each round, which we call a $\beta$-strong adversary, see section 3 for a more formal definition. The adversary can corrupt any message pulled from the nodes under its influence in this round and can arbitrarily change their contents. We do not make any assumptions about the computational power of the adversary. This model covers the special case where the same $\beta n$ nodes are controlled by the adversary throughout all rounds (in other words, the case where there are $\beta n$ adversarial nodes).
@@
With the introduction of such adversarial faults, it is not reasonable to ask the nodes to produce exact answers. Thus, the best that we can hope for is an approximate solution, which is exactly the type of problem we investigate. In the $\varepsilon$-approximate quantile problem, the goal is to compute a value whose rank is in the interval $(\phi-\varepsilon)n,(\phi+\varepsilon)n]$. We will start by tackling the special case $\phi=1/2$ of computing an $\varepsilon$-approximate median, and show afterwards how the problem of computing an approximation of arbitrary quantiles can be reduced to that case. We then turn to the problem of finding the (approximate) mean, which is not as robust to outliers as the median. Thus, computing the mean, even approximately, in the presence of adversarial faults is difficult, as they can sabotage any algorithm by injecting extreme outliers into the computation. The best one can hope for this $\varepsilon$-approximate mean problem is therefore to find a value that is at most an additive $\varepsilon M$ away from the true mean, where the values are assumed to be restricted to the interval $[0,M]$.
@@
The scenario with $\beta n$ adversarial nodes tells us that we cannot hope for an approximation factor $\varepsilon$ smaller than $\beta$. Indeed, one could simply have all the adversarial nodes reliably report an extremely large or an extremely small value, which would shift the quantiles of the “good nodes” by $\beta$, and the same scenario can be applied to the mean approximation problem as well. Hence, all our results have $\varepsilon\geq\beta$ as an (implicit) assumption. In some applications – and given that a $\beta$ fraction of nodes might never output a correct answer anyway – it might be acceptable that some other nodes also hold an incorrect answer, provided this number is small enough, especially if this accelerates the algorithm in return. To investigate the possibility of trading off speed with correctness, we analyze our algorithms from the perspective of parametrized correctness, which is new in this setting. Our gossip algorithms are said to have correctness level $\gamma$ if at most $\gamma n$ nodes are storing an incorrect value (e.g. a value with rank outside $1/2-\varepsilon,1/2+\varepsilon]$ for the $\varepsilon$-approximate median problem) upon termination. Note that we can always assume that $\gamma\geq\frac{1}{2n}$, since $\gamma<\frac{1}{n}$ already implies that there must be no incorrect node.
@@
Paper Organization. The remainder of the paper is organized as follows. We give an overview of our results in section 2, followed by a description of our model and an intuitive description of our three algorithms - complemented by their pseudocode - in section 3. Section 4 contains a brief discussion of related work. Then, we proceed to provide a proof sketch of our main theorems in order: section 5 is dedicated to our median algorithm and the proof of Theorem 1; section 6 concerns itself with our quantile shifting algorithm and the proof of Theorem 2; section 7 deals with our mean algorithm. Then, section 8 contains the discussion of our lower bounds on the round complexity, including a proof of Theorem 4. Finally, in the appendix, we first collect standard concentration results and prove some technical inequalities which are used throughout the paper, then give a complete proof of all our main results.
2 Results
@@
In the remainder of the paper, we say that an event $\mathcal{E}$ occurs with high probability (w.h.p.) if $\mathbb{P}\left[\mathcal{E}\right]=1-n^{-\Omega(1)}$. Our first main result gives a fast and robust algorithm for estimating the median. The pseudocode can be found in Algorithm 1 and is complemented by the runtime and correctness guarantees provided in our first theorem.
@@
We note that the approximation here is weaker as it depends on the promised bound on $M$. The reason for this is simple: the adversary can always claim to have a value that is very large, and this would inevitably skew the computed average to a very large value. Thus, the best one can hope for is to give an error bound as a function of a bound on the possible communicated values.
Note that in the special case where all nodes have values in $\{0,1\}$, the above yields an approximate counting algorithm with an additive error of at most $\varepsilon n$.
@@
We also extend the $\Omega\big{(}\log\log n+\log\frac{1}{\varepsilon}\big{)}$ lower bound given in [HMS18] to an $\Omega\big{(}\log\log\frac{1}{\gamma}+\log\frac{1}{\varepsilon}\big{)}$ lower bound in our setting, see Proposition 5 in section 8. These results prove that our median and quantile algorithms have optimal round complexity and the mean algorithm has almost optimal round complexity. We note that our theorems require varying upper bounds of $\beta$ in terms of $\varepsilon$. As mentioned in the introduction, it is not possible to hope for $\varepsilon$-approximation with $\beta>\varepsilon$, as in particular all adversarial nodes might always output values that are completely out of the range of the correct values. Our analysis here is likely not tight and we did not optimize these bounds since our algorithms already cover a substantial range of values. The same holds for the upper bound on $\varepsilon$ in Theorem 2. For lower bounds of $\varepsilon$, note that any $\varepsilon<\frac{1}{2n}$ would yield an exact algorithm, which is out of reach due to the presence of the adversary. Hence a lower bound which is inversely polynomial in $n$ follows naturally, where again we did not optimize for the precise exponent, and simply stated what was required for our analysis in order to prove concentration of the expected behavior of the algorithms. However, the key reason why we do not view this restriction as major is a different one: For many applications, approximation factors that are a small constant fraction are completely sufficient. Moreover, this is also the parameter range in which our algorithms are extremely fast. More concretely, when both $\varepsilon$ and $\gamma$ are constant, our algorithms are guaranteed to terminate in constant time, no matter how large $n$ is.
3 Model Definition and Algorithm Overview
@@
We consider a network of $n$ nodes that are all connected to each other (one can think of it as a complete graph on $n$ vertices). For all the problems that we study, each node is initially assigned a value that is polynomially bounded in $n$. Communication occurs in synchronous rounds, during each of which every node contacts some node chosen uniformly at random (including itself) and then either pushes or pulls a message - of size $O(\log n)$ bits - from the chosen node. As mentioned before, all our algorithms use the pull operation only, and the message sent out by a pulled node is the same, no matter which (and how many) nodes are pulling from it.
@@
In every round, the messages sent out by at most $\beta n$ of the nodes can become corrupted. More precisely, we assume that these corruptions can be adversarial in the following sense: At the beginning of each round, the adversary can choose at most $\beta n$ nodes so that any message pulled from or pushed to these nodes can be entirely and arbitrarily modified by the adversary, that is, the message can be changed to whatever the adversary wants. Crucially, the adversary does not know beforehand the random choice of partners of the nodes in this round. We do not assume any additional restrictions on the adversary. For example, the adversary may have unbounded computational power and know the entire history of the algorithm at any given moment of time.
@@
Before describing our new adversarially-robust algorithms, we briefly describe the state-of-the-art in the non-adversarial setting. The approximate median find algorithm given by [HMS18] uses a pull-based tournament-style protocol. Here, every node repeatedly updates its value to the median of three randomly chosen node values. The effect is that after $O\big{(}\log\log n+\log\frac{1}{\varepsilon}\big{)}$ rounds, every node has communicated directly or indirectly with $\log n/\varepsilon^{2}$ other nodes. Notice that the problem becomes straightforward with large message sizes, as every node can collect all the information about every node it has directly or indirectly heard from, and $\log n/\varepsilon^{2}$ many uniformly randomly chosen node values are sufficient to prove the concentration of the median value for this node. As the authors show, we do not need to store all such information, and we can make do with just maintaining the median. Their algorithm then concludes with a majority vote over a constant number of nodes to allow for a union bound.
@@
Our algorithms retain the simplicity of these previous algorithms with minor changes while being robust to adversarial corruptions. Some of the crucial changes include giving a more nuanced termination condition which takes into account $\beta$ and $\gamma$, and deriving tailored bounds on expressions involving binomial coefficients to bound the effect of the adversaries on each round. We also introduce an extended second phase – now collecting values from $O\big{(}\frac{\log(1/\gamma)}{\log(1/\beta)}\big{)}$ many nodes – to ensure that at most a $\gamma$ fraction of the nodes are storing an incorrect value upon termination, as desired.
@@
More specifically, for the approximate median computation, as outlined above, we first let the nodes conduct a 3-tournament, that is, in the first phase, each node pulls three node values uniformly at random and updates its own value to the median of the three. In the second, additional, phase, the algorithm then uniformly samples a number of nodes and outputs their median. The pseudocode for our approximate median algorithm is given in Algorithm 1.